# Improving Offline-to-Online Reinforcement Learning with Q-Ensembles

### v3

@@ Abstract
Offline reinforcement learning (RL) is a learning paradigm where an agent learns from a fixed dataset of experience. However, learning solely from a static dataset can limit the performance due to the lack of exploration. To overcome it, offline-to-online RL combines offline pre-training with online fine-tuning, which enables the agent to further refine its policy by interacting with the environment in real-time. Despite its benefits, existing offline-to-online RL methods suffer from performance degradation and slow improvement during the online phase. To tackle these challenges, we propose a novel framework called Ensemble-based Offline-to-Online (E2O) RL. By increasing the number of Q-networks, we seamlessly bridge offline pre-training and online fine-tuning without degrading performance. Moreover, to expedite online performance enhancement, we appropriately loosen the pessimism of Q-value estimation and incorporate ensemble-based exploration mechanisms into our framework. Experimental results demonstrate that E2O can substantially improve the training stability, learning efficiency, and final performance of existing offline RL methods during online fine-tuning on a range of locomotion and navigation tasks, significantly outperforming existing offline-to-online RL methods.
Machine Learning, ICML
[…]
tcb@breakable
@@ 1 Introduction
Reinforcement learning (RL) has shown remarkable success in solving complex decision-making problems, from playing virtual games (Vinyals et al., 2019; Silver et al., 2017, 2018) to controlling tangible robots (Mnih et al., 2015; Tsividis et al., 2021; Schrittwieser et al., 2020). In RL, an agent learns to maximize the cumulative return from large amount of experience data obtained by interacting with an environment. However, in many real-world applications, collecting experience data can be expensive, time-consuming, or even dangerous. This challenge has motivated the development of offline RL, where an agent learns from a fixed dataset of experience collected prior to learning (Fujimoto et al., 2019; Wu et al., 2019).
@@
Offline RL has several advantages over online RL, including the ability to reuse existing data, the potential for faster learning, and the possibility of learning from experiences that are too risky or costly to collect online (Silver et al., 2018). However, offline RL also poses significant challenges, such as the potential for overfiting to the training data and the difficulty of ensuring that the learned policy is safe and optimal in the real-world environment. To address these challenges, offline-to-online RL has emerged as an attractive research direction. This approach combines offline pre-training with online fine-tuning using RL, with the goal of learning from a fixed dataset of offline experience and then continuing to learn online in the real-world environment (Nair et al., 2020; Lee et al., 2022). Offline-to-online RL has the potential to address the limitations of offline RL, such as the sub-optimality of the learned policy. Furthermore, starting with an offline RL policy enables offline-to-online RL to achieve strong performance using fewer online environment samples, compared to collecting large amounts of training data by rolling out policies from scratch.
@@
Prior researches have shown that directly initializing an agent with an offline RL method for online fine-tuning can impede efficient policy improvement due to pessimistic learning (Nair et al., 2020; Zhao et al., 2022). A naive solution to this problem is directly removing the pessimistic term during online training. However, this approach can lead to unstable learning or degraded performance in that the distributional shift between offline datasets and online interactions creates large initial temporal difference errors, causing the oblivion of information learned from offline RL (Lee et al., 2022; Mark et al., 2022). Existing offline-to-online RL methods have attempted to address these challenges through implicit policy constraints (Nair et al., 2020), filtering offline data used for online fine-tuning (Lee et al., 2022; Mao et al., 2022; Mark et al., 2022), adjusting policy constraint weights carefully (Zhao et al., 2022), or training more online policies (Zhang et al., 2023). Nevertheless, these methods still face performance degradation in some tasks and settings, and their performance improvement in the online phase is limited.
Taking inspiration from leveraging Q-ensembles in offline RL (An et al., 2021), we propose a novel approach to address the challenges of offline-to-online RL. Specifically, we conduct comprehensive experiments by discarding the pessimistic term in existing offline RL algorithms and increasing the number of Q-networks in both offline and online phases. We find that Q-ensembles help to alleviate unstable training and performance degradation, and can serve as a more flexible pessimistic term by encompassing various target computation and exploration methods during the online fine-tuning phase. Based on this discovery, we propose an Ensemble-based framework for Offline-to-Online RL (E2O) that bridges offline pre-training and online fine-tuning. We demonstrate the effectiveness of E2O framework by instantiating it on existing offline RL algorithms (Kumar et al., 2020; Chen et al., 2022) across diverse benchmark tasks. The main contributions of this work are summarized as follows:
  • $\bullet$
  • $\bullet$
    We propose a unified framework E2O for offline-to-online RL, which enables a wide range of offline RL algorithms to transition from pessimistic offline pre-training to optimistic online fine-tuning, leading to stable and efficient performance improvement.
  • $\bullet$
    We empirically validate the effectiveness of E2O on various benchmark tasks, including locomotion and navigation tasks, and verify that E2O achieves state-of-the-art performance in comparison to all baseline methods.
2 Related Works
Offline RL
Offline RL algorithms focus on training RL agents with pre-collected datasets. However, these algorithms face the challenge of distribution shift between the behavior policy and the policy being learned, which can cause issues due to out-of-distribution (OOD) actions sampled from the learned policy and passed into the learned critic. To mitigate this problem, prior methods constrain the learned policy to stay close to the behavior policy via explicit policy regularization (Fujimoto et al., 2019; Kumar et al., 2019; Wu et al., 2019), via implicit policy constraints (Peng et al., 2019; Siegel et al., 2020; Kostrikov et al., 2021; Zhou et al., 2021; Chen et al., 2022), by leveraging auxiliary behavioral cloning losses (Fujimoto & Gu, 2021), by penalizing the Q-value of OOD actions to prevent selecting them (Kumar et al., 2020; Lyu et al., 2022; An et al., 2021; Bai et al., 2022; Yang et al., 2022), or through model-based training with conservative penalties (Yu et al., 2020; Kidambi et al., 2020; Argenson & Dulac-Arnold, 2020; Yu et al., 2021). Among the above methods, we choose several representative algorithms such as CQL (Kumar et al., 2020) and LAPO (Chen et al., 2022) to be the base methods in the offline component of our framework due to their wide applicability and superior performance.
Offline-to-Online RL
Offline-to-online RL refers to the process of improving the well-trained offline policy by incorporating online interactions. Directly applying pre-trained offline policy to the online fine-tuning stage may lead to poor performance due to excess conservatism (Nair et al., 2020; Lee et al., 2022; Zhao et al., 2022). To adapt offline RL algorithms to the online environment, several modifications are required. AWAC (Nair et al., 2020) is the first algorithm proposed to perform well in the offline-to-online RL setting, which forces the policy to imitate actions with high advantage estimates. AW-Opt (Lu et al., 2021) improves upon AWAC by incorporating positive sample filtering and hybrid actor-critic exploration during the online stage. Balanced Replay (Lee et al., 2022) trains an additional neural network to prioritize samples in order to effectively use new data as well as near-on-policy samples from the offline dataset. PEX (Zhang et al., 2023) proposes a policy expansion approach for offline-to-online RL, which trains more policies from scratch in the online phase and combine them with the pre-trained offline policy to make decisions jointly. A concurrent work with us is O3F (Mark et al., 2022), which also aims to eliminate pessimism in online learning. However, our research takes a more holistic perspective by connecting the pessimistic offline learning and optimistic online phases through ensemble modeling, proposing a more applicable framework that encapsulates the implementation of O3F as a special case.
Q-Ensembles in RL
Q-Ensemble methods have been widely utilized to enhance the performance of RL (Van Hasselt et al., 2016; Osband et al., 2016; Anschel et al., 2017; Fujimoto et al., 2018; Ciosek et al., 2019; Lan et al., 2020; Chen et al., 2021; Lee et al., 2021). TD3 (Fujimoto et al., 2018) leverages an ensemble of two value functions and uses their minimum for computing the target value during Bellman error minimization. REDQ (Chen et al., 2021) minimizes over a random subset of Q-functions in the target to reduce over-estimation bias. For offline RL, a number of works have extended this to propose backing up minimums or lower confidence bound estimates over larger ensembles (Fujimoto et al., 2019; Kumar et al., 2019; Wu et al., 2019; Agarwal et al., 2020; An et al., 2021; Bai et al., 2022; Yang et al., 2022). In particular, EDAC (An et al., 2021) achieves impressive performance by simply increasing the number of Q-networks along with the clipped Q-learning and further proposes to reduce the required number of ensemble networks through ensemble gradient diversification for the purpose of reducing computational cost. PBRL (Bai et al., 2022) and RORL (Yang et al., 2022) both employ an ensemble of bootstrapped Q-functions for uncertainty quantification and perform pessimistic updates to penalize Q functions with high uncertainties. Recent work (Ghasemipour et al., 2022) advocates for using independently learned ensembles without sharing target values and optimizing a policy based on the lower confidence bound of predicted action values. In our work, we discover that Q-ensembles are highly effective in addressing the performance degradation that occurs in the offline-to-online setting, and we advocate for using Q-ensembles for both offline and online algorithms to achieve steady and sample-efficient offline-to-online RL.
Exploration Mechanisms
Various exploration approaches have been proposed to accelerate the efficiency of online training in recent years. These methods can typically be divided into two main categories following (Hao et al., 2023): uncertainty-oriented exploration and intrinsic motivation-oriented exploration. The former employs heuristic design to formulate various intrinsic motivations for exploration, based on factors such as visitation count (Bellemare et al., 2016; Ostrovski et al., 2017; Tang et al., 2017), curiosity (Savinov et al., 2018), information gain (Houthooft et al., 2016), etc. Uncertainty-oriented exploration methods adopt the principle of optimism in the face of uncertainty and use Q-ensembles to encourage agents to explore areas with higher uncertainty (Badia et al., 2020; Ecoffet et al., 2021; Lee et al., 2021; Ishfaq et al., 2021). In this paper, we investigate uncertainty-based exploration mechanisms under the unified framework of Q-ensembles to enhance the performance during the online fine-tuning phase.
3 Why Can Q-Ensembles Help Offline-to-Online RL?
@@
To get a better understanding of our ensemble-based framework, we begin with examples that highlight the advantages of Q-ensembles for offline-to-online RL. A natural starting point for offline-to-online RL is to simply initialize the agent with the one trained by an existing offline RL method and then directly perform online fine-tuning without using the offline dataset. However, this approach can hinder efficient online performance improvement due to the inherent pessimism of the offline learning paradigm (Lee et al., 2022; Mark et al., 2022). To support this claim, we present CQL (Kumar et al., 2020) as a representative and conduct preliminary experiments on the D4RL Walker2d-medium-expert-v2 dataset. The learning curve of CQL during online fine-tuning in Figure 1(a) shows that CQL can maintain the offline performance at the initial stage of online fine-tuning and steadily improve during the training process. This can be attributed to the use of pessimistic Q-functions, which enables the agent to visit states resembling those in the offline dataset and maintain pessimistic towards unseen actions during the initial stage of online fine-tuning. However, the pessimistic objective impedes proper exploration in the online stage and restrict the agent from efficiently improving its performance (Lee et al., 2022; Mark et al., 2022).
@@
To tackle the aforementioned issue of limited exploration, one might be tempted to remove the conservative estimation component in order to reduce the conservatism of the learning process. However, as shown in Figure 1(a), this naive solution leads to unstable training or performance degradation when switching from CQL to Soft Actor-Critic (SAC) (Haarnoja et al., 2018) during online fine-tuning, which has also been reported in previous offline-to-online RL works (Lu et al., 2021; Nair et al., 2020; Lee et al., 2022; Mark et al., 2022). The reason is that SAC lacks accurate estimation of Q-values for unknown state-action pairs. Without the conservative constraints of CQL, the Q-values tend to be overestimated, leading to policy misguidance.
@@
So is it possible to find a method that retains suitable pessimistic constraints to mitigate performance degradation, while also tailoring these constraints to be more conducive to exploration during the online phase, rather than being as conservative as traditional offline RL algorithms such as CQL? Inspired by increasing the number of Q-networks in (An et al., 2021), we introduce Q-ensembles and set the number of Q functions in CQL and SAC to N. Specifically, the target Q value is estimated by selecting the minimum value from all the Q-ensembles. We refer to these intermediate methods as CQL-N and SAC-N. Figure 1(a) shows the effectiveness of using SAC-N for online fine-tuning of an offline policy pre-trained with CQL-N. Surprisingly, after incorporating Q-ensembles, we observe that the training becomes more stable and performance drop is no longer observed when switching to online fine-tuning. Moreover, this constraint method not only enhances the final performance of the offline stage, but also improves the efficiency of online learning.
@@
We have seen that both SAC-N and CQL can prevent performance drop during online fine-tuning, but why does SAC-N exhibit better performance compared to CQL? To answer this question, we analyze the distance between the actions selected by each method and the actions in the dataset, as shown in Figure 1(c). Specifically, we measure $\mathbb{E}_{(\mathbf{s},\mathbf{a})\sim\mathcal{D},\hat{\mathbf{a}}\sim\pi_{\theta}(\cdot\mid\mathbf{s})}\left[\|\hat{\mathbf{a}}-\mathbf{a}\|_{2}^{2}\right]$ for SAC-N, CQL and a random policy by performing online fine-tuning on the Walker2d-medium-expert-v2 dataset. Our findings reveal that SAC-N has a wider range of action choices compared to CQL, and a more diverse set of actions can lead to improved performance, as stated in previous exploration methods (Badia et al., 2020; Ecoffet et al., 2021; Lee et al., 2021). Therefore, we can incorporate Q-ensembles into existing offline RL algorithms like CQL, and discard the original conservative term designed for offline algorithms during the online phase to improve the online learning efficiency.
@@ 4 Ensemble-based Offline-to-Online Reinforcement Learning
Based on the empirical observations discussed earlier, we propose our Ensemble-based Offline-to-Online (E2O) RL Framework. In this section, we first present merits of Q-ensemble using additional empirical results and then progressively introduce more ensemble-based mechanisms into our framework. Although each individual design decision in E2O may seem relatively simple, their specific combination outperforms baselines in terms of training stability, learning efficiency and final performance.
@@ 4.1 Q-Ensembles
As discussed in Section 3, Q-ensembles can bridge offline and online phases to help pre-trained offline agents perform stable online fine-tuning. In this section, we present comprehensive empirical results to further verify its advantages.
Given an offline RL algorithm named OfflineRL, we introduce Q-ensembles to get OfflineRL-N, indicating that the algorithm uses $N$ Q-networks and takes the minimum value of all the Q-networks in the ensemble to compute the target. With the pre-trained OfflineRL-N agent, we load it as the initialization of the online agent and remove the originally designed pessimistic term (if possible) to obtain OnlineRL-N. Then OnlineRL-N is trained online. In all methodology sections, we instantiate OfflineRL as CQL, and thus OfflineRL-N refers to CQL-N, and OnlineRL-N refers to SAC-N, as described in Section 3. To comprehensively verify the effectiveness of Q-ensembles in stabilizing training process and mitigating performance drop, we consider three MuJoCo locomotion tasks (Todorov et al., 2012): HalfCheetah, Hopper, and Walker2d from the D4RL benchmark suite (Fu et al., 2020). Specifically, we consider the medium, medium-replay and medium-expert datasets, as in typical real-world scenarios, we rarely use a random policy or have an expert policy for system control.
@@
Figure 2 shows the aggregated normalized return across all nine datasets. Consistent with the results of the previous illustrative experiment, online training of OfflineRL is stable but leads to slower asymptotic performance. Directly switching to OnlineRL causes unstable training process and performance drop. In contrast, OfflineRL-N normal-$\to$ OnlineRL-N no longer experiences performance collapse after switching to online fine-tuning, and the training process is relatively stable. Additionally, OfflineRL-N normal-$\to$ OnlineRL-N achieves better fine-tuned performance than OfflineRL normal-$\to$ OfflineRL.
Although the ensemble-based method OfflineRL-N normal-$\to$ OnlineRL-N has made certain improvements compared to existing method OfflineRL normal-$\to$ OfflineRL, it still fails to be improved rapidly in the online stage compared with standard online RL algorithms. Therefore, we shift our focus to analyzing whether we can appropriately loosen the pessimistic estimation of Q-values in the online phase to further improve learning efficiency while ensuring stable training.

### v4

# ENOTO: Improving Offline-to-Online Reinforcement Learning with Q-Ensembles

@@ Abstract
Offline reinforcement learning (RL) is a learning paradigm where an agent learns from a fixed dataset of experience. However, learning solely from a static dataset can limit the performance due to the lack of exploration. To overcome it, offline-to-online RL combines offline pre-training with online fine-tuning, which enables the agent to further refine its policy by interacting with the environment in real-time. Despite its benefits, existing offline-to-online RL methods suffer from performance degradation and slow improvement during the online phase. To tackle these challenges, we propose a novel framework called ENsemble-based Offline-To-Online (ENOTO) RL. By increasing the number of Q-networks, we seamlessly bridge offline pre-training and online fine-tuning without degrading performance. Moreover, to expedite online performance enhancement, we appropriately loosen the pessimism of Q-value estimation and incorporate ensemble-based exploration mechanisms into our framework. Experimental results demonstrate that ENOTO can substantially improve the training stability, learning efficiency, and final performance of existing offline RL methods during online fine-tuning on a range of locomotion and navigation tasks, significantly outperforming existing offline-to-online RL methods.
@@ 1 Introduction
Reinforcement learning (RL) has shown remarkable success in solving complex decision-making problems, from playing virtual games Silver et al. (2017); Vinyals et al. (2019) to controlling tangible robots Mnih et al. (2015); Tsividis et al. (2021); Schrittwieser et al. (2020). In RL, an agent learns to maximize the cumulative return from large amount of experience data obtained by interacting with an environment. However, in many real-world applications, collecting experience data can be expensive, time-consuming, or even dangerous. This challenge has motivated the development of offline RL, where an agent learns from a fixed dataset of experience collected prior to learning Fujimoto et al. (2019); Wu et al. (2019); Bai et al. (2022); Liu et al. (2023); Yu et al. (2020); Kidambi et al. (2020).
@@
Offline RL has several advantages over online RL, including the ability to reuse existing data, the potential for faster learning, and the possibility of learning from experiences that are too risky or costly to collect online Silver et al. (2018). However, offline RL also poses significant challenges, such as the potential for overfiting to the training data and the difficulty of ensuring that the learned policy is safe and optimal in the real-world environment. To address these challenges, offline-to-online RL has emerged as an attractive research direction. This approach combines offline pre-training with online fine-tuning using RL, with the goal of learning from a fixed dataset of offline experience and then continuing to learn online in the real-world environment Nair et al. (2020); Lee et al. (2022). Offline-to-online RL has the potential to address the limitations of offline RL, such as the sub-optimality of learned policy. Furthermore, starting with an offline RL policy can achieve strong performance with fewer online environment samples, compared to collecting large amounts of training data by rolling out policies from scratch.
@@
Prior researches have shown that directly initializing an agent with an offline RL method for online fine-tuning can impede efficient policy improvement due to pessimistic learning Nair et al. (2020); Zhao et al. (2022). A naive solution to this problem is directly removing the pessimistic term during online training. However, this approach can lead to unstable learning or degraded performance in that the distributional shift between offline datasets and online interactions creates large initial temporal difference errors, causing the oblivion of information learned from offline RL Lee et al. (2022); Mark et al. (2022). Existing offline-to-online RL methods have attempted to address these challenges through implicit policy constraints Nair et al. (2020), filtering offline data used for online fine-tuning Lee et al. (2022); Mao et al. (2022); Mark et al. (2022), adjusting policy constraint weights carefully Zhao et al. (2022), or training more online policies Zhang et al. (2023). Nevertheless, these methods still face performance degradation in some tasks and settings, and their performance improvement in the online phase is limited.
@@
Taking inspiration from leveraging Q-ensembles in offline RL An et al. (2021), we propose a novel approach to address the challenges of offline-to-online RL. Specifically, we conduct comprehensive experiments by discarding the pessimistic term in existing offline RL algorithms and increasing the number of Q-networks in both offline and online phases. We find that Q-ensembles help to alleviate unstable training and performance degradation, and can serve as a more flexible pessimistic term by encompassing various target computation and exploration methods during the online fine-tuning phase. Based on this discovery, we propose an ENsemble-based Offline-To-Online (ENOTO) RL framework that bridges offline pre-training and online fine-tuning. We demonstrate the effectiveness of ENOTO framework by instantiating it on existing offline RL algorithms Kumar et al. (2020); Chen et al. (2022) across diverse benchmark tasks. The main contributions of this work are summarized as follows:
  • We propose a unified framework ENOTO for offline-to-online RL, which enables a wide range of offline RL algorithms to transition from pessimistic offline pre-training to optimistic online fine-tuning, leading to stable and efficient performance improvement.
  • We empirically validate the effectiveness of ENOTO on various benchmark tasks, including locomotion and navigation tasks, and verify that ENOTO achieves state-of-the-art performance in comparison to all baseline methods.
2 Why Can Q-Ensembles Help Offline-to-Online RL?
@@
To get a better understanding of our ensemble-based framework, we begin with examples that highlight the advantages of Q-ensembles for offline-to-online RL. A natural starting point for offline-to-online RL is to simply initialize the agent with the one trained by an existing offline RL method and then directly perform online fine-tuning without using the offline dataset. However, this approach can hinder efficient online performance improvement due to the inherent pessimism of the offline learning paradigm Lee et al. (2022); Mark et al. (2022). To support this claim, we present CQL Kumar et al. (2020) as a representative and conduct preliminary experiments on the D4RL Walker2d-medium-expert-v2 dataset. The learning curve of CQL during online fine-tuning in Fig. 1(a) shows that CQL can maintain the offline performance at the initial stage of online fine-tuning and steadily improve during the training process. This can be attributed to the use of pessimistic Q-functions, which enables the agent to visit states resembling those in the offline dataset and maintain pessimistic towards unseen actions during the initial stage of online fine-tuning. However, the pessimistic objective impedes proper exploration in the online stage and restrict the agent from efficiently improving its performance Lee et al. (2022); Mark et al. (2022); Hao et al. (2023); Ghasemipour et al. (2022).
@@
To tackle the aforementioned issue of limited exploration, one might be tempted to remove the conservative estimation component in order to reduce the conservatism of the learning process. However, as shown in Fig. 1(a), this naive solution leads to unstable training or performance degradation when switching from CQL to Soft Actor-Critic (SAC) Haarnoja et al. (2018) during online fine-tuning, which has also been reported in previous offline-to-online RL works Lu et al. (2021); Nair et al. (2020); Lee et al. (2022); Mark et al. (2022). The reason is that SAC lacks accurate estimation of Q-values for unknown state-action pairs. Without the conservative constraints of CQL, the Q-values tend to be overestimated, leading to policy misguidance.
@@
So is it possible to find a method that retains suitable pessimistic constraints to mitigate performance degradation, while also tailoring these constraints to be more conducive to exploration during the online phase, rather than being as conservative as traditional offline RL algorithms such as CQL? Inspired by increasing the number of Q-networks in An et al. (2021), we introduce Q-ensembles and set the number of Q functions in CQL and SAC to N. Specifically, the target Q value is estimated by selecting the minimum value from all the Q-ensembles. We refer to these intermediate methods as CQL-N and SAC-N. Fig. 1(a) shows the effectiveness of using SAC-N for online fine-tuning of an offline policy pre-trained with CQL-N. Surprisingly, after incorporating Q-ensembles, we observe that the training becomes more stable and performance drop is no longer observed when switching to online fine-tuning. Moreover, this constraint method not only enhances the final performance of the offline stage, but also improves the efficiency of online learning.
@@
We have seen that both SAC-N and CQL can prevent performance drop during online fine-tuning, but why does SAC-N exhibit better performance compared to CQL? To answer this question, we analyze the distance between the actions selected by each method and the actions in the dataset, as shown in Fig. 1(c). Specifically, we measure for SAC-N, CQL and a random policy by performing online fine-tuning on the Walker2d-medium-expert-v2 dataset. Our findings reveal that SAC-N has a wider range of action choices compared to CQL, and a more diverse set of actions can lead to improved performance, as stated in previous exploration methods Ecoffet et al. (2021); Lee et al. (2021); Liu et al. (2024); Savinov et al. (2018); Houthooft et al. (2016). Therefore, we can incorporate Q-ensembles into existing offline RL algorithms like CQL, and discard the original conservative term designed for offline algorithms during the online phase to improve the online learning efficiency.
@@ 3 Ensemble-based Offline-to-Online Reinforcement Learning
Based on the empirical observations discussed earlier, we propose our ENnsemble-based Offline-To-Online (ENOTO) RL Framework. In this section, we first present merits of Q-ensemble using additional empirical results and then progressively introduce more ensemble-based mechanisms into our framework. Although each individual design decision in ENOTO may seem relatively simple, their specific combination outperforms baselines in terms of training stability, learning efficiency and final performance.
@@ 3.1 Q-Ensembles
As discussed in the previous section, Q-ensembles can bridge offline and online phases to help pre-trained offline agents perform stable online fine-tuning. In this section, we present comprehensive empirical results to further verify its advantages.
Given an offline RL algorithm named OfflineRL, we introduce Q-ensembles to get OfflineRL-N, indicating that the algorithm uses $N$ Q-networks and takes the minimum value of all the Q-networks in the ensemble to compute the target. With the pre-trained OfflineRL-N agent, we load it as the initialization of the online agent and remove the originally designed pessimistic term (if possible) to obtain OnlineRL-N. Then OnlineRL-N is trained online. In all methodology sections, we instantiate OfflineRL as CQL, and thus OfflineRL-N refers to CQL-N, and OnlineRL-N refers to SAC-N. To comprehensively verify the effectiveness of Q-ensembles in stabilizing training process and mitigating performance drop, we consider three MuJoCo locomotion tasks Todorov et al. (2012): HalfCheetah, Hopper, and Walker2d from the D4RL benchmark suite Fu et al. (2020). Specifically, we consider the medium, medium-replay and medium-expert datasets, as in typical real-world scenarios, we rarely use a random policy or have an expert policy for system control.
Fig. 2 shows the aggregated normalized return across all nine datasets. Consistent with the results of the previous illustrative experiment, online training of OfflineRL is stable but leads to slower asymptotic performance. Directly switching to OnlineRL causes unstable training process and performance drop. In contrast, OfflineRL-N $\to$ OnlineRL-N no longer experiences performance collapse after switching to online fine-tuning, and the training process is relatively stable. Additionally, OfflineRL-N $\to$ OnlineRL-N achieves better fine-tuned performance than OfflineRL $\to$ OfflineRL.
Although the ensemble-based method OfflineRL-N $\to$ OnlineRL-N has made certain improvements compared to existing method OfflineRL $\to$ OfflineRL, it still fails to be improved rapidly in the online stage compared with standard online RL algorithms. Therefore, we shift our focus to analyzing whether we can appropriately loosen the pessimistic estimation of Q-values in the online phase to further improve learning efficiency while ensuring stable training.